# SHMC-NET: A Mask-guided Feature Fusion Network for Sperm Head Morphology Classification

### 06 Feb 2024 v1

@@ Abstract
Male infertility accounts for about one-third of global infertility cases. Manual assessment of sperm abnormalities through head morphology analysis encounters issues of observer variability and diagnostic discrepancies among experts. Its alternative, Computer-Assisted Semen Analysis (CASA), suffers from low-quality sperm images, small datasets, and noisy class labels. We propose a new approach for sperm head morphology classification, called SHMC-Net, which uses segmentation masks of sperm heads to guide morphology classification of sperm images. SHMC-Net generates reliable segmentation masks using image priors, refines object boundaries with an efficient graph-based method, and trains an image network with sperm head crops and a mask network with the corresponding masks. In the intermediate stages of the networks, image and mask features are fused with a fusion scheme to better learn morphological features. To handle noisy class labels and regularize training on small datasets, SHMC-Net applies Soft Mixup to combine mixup augmentation and a loss function. We achieve state-of-the-art results on SCIAN and HuSHeM datasets, outperforming methods that use additional pre-training or costly ensembling techniques.
@@ 1 Introduction
To address these challenges we propose to incorporate additional information from sperm head masks to help morphology classification. Sperm head masks can be reliably and efficiently extracted from images, as shown in previous work [17, 3, 24, 8]. Such masks in general have well-delineated boundaries with fewer background artifacts. The objects of interest in the masks are represented as the foreground with morphologically relevant shape, size, and structure information that is easy to extract while containing little distracted structures that may inhibit shape analysis. Motivated by these observations, we propose a new approach for effective Sperm Head Morphology Classification, called SHMC-Net. Our major contributions are (1) a new sperm head morphology-aware classification approach that explicitly utilizes morphology information from sperm head segmentation masks, (2) a classification network with a novel architecture, Fusion Encoder, that applies an effective feature fusion scheme to utilize information from both the input images and masks (3) a novel combination of augmentation and loss function to handle noisy labels and regularize training on small datasets, (4) an efficient graph-based sperm head mask boundary refinement method and (5) SOTA performances on 2 public sperm datasets outperforming even the best-known methods that use additional pre-training or costly ensembling of multiple convolution neural network (CNN) models.
@@ 2.1 Mask Generation and Refinement
Reasonable head masks can be extracted from sperm images [17, 3, 24, 8] in an unsupervised fashion but these masks may have imperfect boundaries. Hence, given a raw sperm image, we first generate segmentation masks for sperm head crops using anatomical and image priors and further refine mask boundaries using a graph-based method. Using the HPM method [24], we obtain the sperm-head-only crops from the sperm image, right-align them, and produce their initial pseudo-masks. We propose a sperm head shape-aware Graph-based Boundary Refinement (GrBR) method. GrBR generates the boundary-refined masks instantaneously ($<7$ ms per image). The GrBR module refines the boundary contour $C$ of the HPM pseudo-mask in a head crop to obtain a closed curve $C^{\prime}$ that captures the accurate head boundary of the mask. We formulate the optimal contour refinement problem as computing the shortest path in a directed graph $G$ defined on $C$.
We first uniformly sample $n$ points clockwise on $C$, and for each sample point $a_{j}\in C$, a line segment $l(a_{j})$ is drawn orthogonal to the curvature of $C$ at $a_{j}$, with $a_{j}$ being the center point of $l(a_{j})$. We then sample $m$ points uniformly on each line segment $l(a_{j})$. Let $p_{i}^{j}:=(x(p_{i}^{j}),y(p_{i}^{j}))$ represent the $i$-th sample point on the line segment $l(a_{j})$, $i=1,2,\ldots,m$. All the sample points on $C$ and on the line segments are vertices of the graph $G$ and have weights equal to the negatives of their gradients. Since sperm heads have almost convex shapes, we impose a smoothness constraint and a near convex constraint for computing $C^{\prime}$. The smoothness constraint is enforced by using a smoothness parameter $s$ such that each $p_{i}^{j}$ can have a directed edge to any point $p_{i^{\prime}}^{j+1-\lfloor j/n\rfloor\times n}$ only, for each $i^{\prime}$ with $|i^{\prime}-i|\leq s$, where $i=1,2,\dots,m$, $i^{\prime}=1,2,\dots,m$, and $j=1,2,\dots,n$. The shape constraint is enforced by penalizing any concave edge-to-edge connection along $C^{\prime}$ with a larger connection cost, i.e., edge weight = $c\times(\theta-\pi)$ if $\theta>\pi$, and else edge weight = 0, where $c$ = cost penalty = initial edge weight and $\theta$ is the angle between the two adjacent edges. We make $G$ into a directed acyclic graph by making a copy of all the vertices on the first line segment $l(a_{1})$. These vertex copies are now on the line segment copy $l^{\prime}(a_{1})$ and their edges are from the vertices of $l(a_{n})$.
An optimal closed curve $C^{\prime}$ in the head crop subject to the smoothness and shape constraints based on $C$ is obtained by computing the shortest path in the graph $G$ in polynomial time using dynamic programming. The obtained $C^{\prime}$ corresponds to the boundary-refined sperm head mask.
@@ 2.2 Fusion Encoder
We propose a novel Fusion Encoder, to leverage the mask features and facilitate the morphology classification. Using the masks directly to train a classification network is not ideal (see Table 4) because the masks lack texture details and are polarized to all-or-nothing errors due to their coarse discretized representation. Hence, we use a one-directional feature fusion scheme (from mask network to image), at the deeper network layers (evidence for this design in Section 3.5) to exploit the learned semantics from both the image and mask networks. The Fusion Encoder consists of an image network, a mask network, fusion blocks, and a classifier.
2.2.1 Image and Mask Network
We use two 5-staged lightweight ShuffleNet [12] backboned encoders as our image and mask network. In the first three stages, both networks learn the features independent from each other. After the third stage, features from the image and mask networks are fused and passed to the fourth stage of the image network. The features from the two networks are fused again before the fifth stage and passed to the image network to obtain the final features from the image network.
2.2.2 Fusion Blocks and Classifier
Fusion blocks are placed at the beginning of the $4^{th}$ and $5^{th}$ stages of the networks and consist of a pixel-wise summation of the image and mask features, followed by a convolution, batch normalization, and ReLU operations. A pixel-wise summation is performed between the final features obtained at the end of the image and mask networks, followed by average pooling and fed through the classifier. We use a randomly initialized fully connected linear layer as a classifier that outputs (after the softmax and argmax operations) the morphology class predictions of the input sperm head crop.
@@ 2.3 Soft Mixup Regularization
Accurately determining the sperm head morphology classes is difficult even for experts, which may lead to disagreement on expert-labeled class IDs of samples (i.e., for the SCIAN dataset a sample may have a majority class label with consensus among at least 2/3 expert labels and a minority class label with no consensus).
2.3.1 Soft Mixup Augmentation
We apply an intra-class Mixup augmentation [23] scheme to handle samples with disagreed labels and reduce class conditional input variance. Given a sample image crop $x_{i}$ and its corresponding mask $m_{i}$ with a majority label $y^{a}_{i}$ and a minority label $y^{b}_{i}$, we randomly pick another sample $x_{j}$ (with a mask $m_{j}$, a majority label $y^{a}_{j}$, and a minority label $y^{b}_{j}$) if $y^{a}_{i}=y^{a}_{j}$. Prior to this step, we perform random oversampling of all the under-represented classes to match the number of samples in the most represented class. We then apply separate image augmentation ($\Phi$) to $x_{i}$ and $x_{j}$. The input, $\ddot{x}_{i}$, to the Fusion Encoder is a linear combination of the augmented mixup samples $\Phi_{i}(x_{i})$ and $\Phi_{j}(x_{j})$, and its target label $\ddot{y}_{i}$ is equivalent to the majority label ($y_{i}^{a}$). Formally:
@@ 2.3.1 Soft Mixup Augmentation
2.3.2 Soft Mixup Loss
We modify the Soft Loss in [24] to adapt intra-class mixup to handle partial agreements among the morphology class labels. Soft Loss is defined as:
@@ 3.1 Datasets
We demonstrate the performance of SHMC-Net on two public sperm morphology datasets. SCIAN-Morpho Sperm GS (SCIAN) [1] contains 1854 gray-scale sperm images of size 35 $\times$ 35 (height x width), each belonging to one of five morphological classes: Normal (N), Tapered (T), Pyriform (P), Small (S), or Amorphous (A) as labeled by 3 experts. Samples where all 3 experts agree on the same class label form the SCIAN total Agreement (SCIAN-TA) dataset (384 images: 35 N, 69 T, 7 P, 11 S, 262 A), and the samples where at least 2 experts agree on class labels form the SCIAN Partial Agreement (SCIAN-PA) dataset (1132 images: 100 N, 228 T, 76 P, 72 S, 656 A)
@@ 3.2 Implementation Details
ImageNet pre-trained ShuffleNet V2 [12] is used as the backbone for our image and mask networks and trained using the AdamW optimizer with an initial learning rate of 0.00015 and the Cosine Annealing learning rate scheduler. The model is trained on a single NVIDIA TITAN-Xp GPU for 300 epochs (batch size of 128 for SCIAN-PA and 64 for SCIAN-TA and HuSHeM). Dropout with 40% probability and weight decay of 0.01 was applied as additional regularization. The majority class weight $\gamma$ is set to 0.85 and mixup strength, $\lambda$, is sampled randomly from the beta distribution with $\alpha=0.5$. All the image crops are first resized to 64$\times$64 and randomly square-cropped between sizes 45 to 64 and resized again to 64*64 along with 10-degree rotation, vertical flip, and horizontal/vertical shift by up to 10% of their width or height. We evaluate the performance in terms of Overall Accuracy, macro-averaged class-wise Precision and Recall, and the F1 score attained as a harmonic mean of the two. We apply 5-fold cross-validation and report the average of 5 runs along with the standard deviation.
@@ 3.5 Component Studies
Method Ensemble
Accuracy
Precision Recall F1
@@
Spencer et al. [21]
98.52
98.53 98.52 98.52
SHMC-Net-E (ours)
99.17 $\pm$ 0.20
99.19 $\pm$ 0.21 99.16 $\pm$ 0.20 99.17 $\pm$ 0.20
@@ 3.6 Ablation Study
Using only masks to train the model yields poor performance (Table 4) since their coarse discretized representation lacks texture details and is polarized to all-or-nothing errors. However, combining mask and image features at the classifier level has improvements (3.4% F1 on SCIAN-PA and 1.8% on HuSHeM) over using image-only. Intermediate stage fusion of mask and image features further strengthens the model’s ability to learn morphology features (further 1.9% on SCIAN-PA and 1.3% on HuSHeM). Finally, handling class imbalance and label discrepancies through Soft Mixup adds 2.2% further F1 improvement on SCIAN-PA, and the regularization through Soft Mixup contributes to an additional 0.4% F1 on the small-size HuSHeM dataset.

### 05 Mar 2024 v3

@@ Abstract
Male infertility accounts for about one-third of global infertility cases. Manual assessment of sperm abnormalities through head morphology analysis encounters issues of observer variability and diagnostic discrepancies among experts. Its alternative, Computer-Assisted Semen Analysis (CASA), suffers from low-quality sperm images, small datasets, and noisy class labels. We propose a new approach for sperm head morphology classification, called SHMC-Net, which uses segmentation masks of sperm heads to guide morphology classification of sperm images. SHMC-Net generates reliable segmentation masks using image priors, refines object boundaries with an efficient graph-based method, and trains an image network with sperm head crops and a mask network with the corresponding masks. In the intermediate stages of the networks, image and mask features are fused with a fusion scheme to better learn morphological features. To handle noisy class labels and regularize training on small datasets, SHMC-Net applies Soft Mixup to combine mixup augmentation and a loss function. We achieve state-of-the-art results on SCIAN and HuSHeM datasets. Code is available in GitHub.
@@ 1 Introduction
To address these challenges we propose to incorporate additional information from sperm head masks to help morphology classification. Sperm head masks can be reliably and efficiently extracted from images, as shown in previous work [15, 3, 21, 8]. Such masks in general have well-delineated boundaries with fewer background artifacts. The objects of interest in the masks are represented as the foreground with morphologically relevant shape, size, and structure information that is easy to extract while containing little distracted structures that may inhibit shape analysis. Motivated by these observations, we propose a new approach for effective Sperm Head Morphology Classification, called SHMC-Net. Our major contributions are (1) a new sperm head morphology-aware classification network architecture, Fusion Encoder, that applies an effective feature fusion scheme to utilize morphology information from both the input images and segmentation masks (3) a novel combination of augmentation and loss function to handle noisy labels and regularize training on small datasets, (4) an efficient graph-based sperm head mask boundary refinement method and (5) SOTA performances on 2 public sperm datasets outperforming even the best-known methods that use additional pre-training or costly ensembling of multiple convolution neural network (CNN) models.
@@ 2.1 Mask Generation and Refinement
Reasonable head masks can be extracted from sperm images in an unsupervised fashion [15, 3, 21, 8] but these masks may have imperfect boundaries. Hence, given a raw sperm image, we first generate segmentation masks for sperm head crops using anatomical and image priors and further refine mask boundaries. Using the HPM method [21], we obtain the sperm-head-only crops from the sperm image, right-align them, and produce their pseudo-masks. We propose a sperm head shape-aware Graph-based Boundary Refinement (GrBR) method. GrBR generates the boundary-refined masks instantaneously ($<7$ ms per image). The GrBR module refines the boundary contour $C$ of the HPM pseudo-mask in a head crop to obtain a closed curve $C^{\prime}$ that captures the accurate head boundary of the mask. We formulate the optimal contour refinement problem as computing the shortest path in a directed graph $G$ defined on $C$. We first uniformly sample $n$ points clockwise on $C$, and for each sample point $a_{j}\in C$, a line segment $l(a_{j})$ is drawn orthogonal to the curvature of $C$ at $a_{j}$, with $a_{j}$ being the center point of $l(a_{j})$. We then sample $m$ points uniformly on each line segment $l(a_{j})$. Let $p_{i}^{j}:=(x(p_{i}^{j}),y(p_{i}^{j}))$ represent the $i$-th sample point on the line segment $l(a_{j})$, $i=1,2,\ldots,m$. All the sample points on $C$ and on the line segments are vertices of the graph $G$ and have weights equal to the negatives of their gradients. Since sperm heads have almost convex shapes, we impose a smoothness constraint and a near convex constraint for computing $C^{\prime}$. The smoothness constraint is enforced by using a smoothness parameter $s$ such that each $p_{i}^{j}$ can have a directed edge to any point $p_{i^{\prime}}^{j+1-\lfloor j/n\rfloor\times n}$ only, for each $i^{\prime}$ with $|i^{\prime}-i|\leq s$, where $i=1,2,\dots,m$, $i^{\prime}=1,2,\dots,m$, and $j=1,2,\dots,n$. The shape constraint is enforced by penalizing any concave edge-to-edge connection along $C^{\prime}$ with a larger connection cost, i.e., edge weight = $c\times(\theta-\pi)$ if $\theta>\pi$, and else edge weight = 0, where $c$ = cost penalty = initial edge weight and $\theta$ is the angle between the two adjacent edges. We make $G$ into a directed acyclic graph by making a copy of all the vertices on the first line segment $l(a_{1})$. These vertex copies are now on the line segment copy $l^{\prime}(a_{1})$ and their edges are from the vertices of $l(a_{n})$. An optimal closed curve $C^{\prime}$ in the head crop subject to the smoothness and shape constraints based on $C$ is obtained by computing the shortest path in the graph $G$ in polynomial time using dynamic programming.
@@ 2.2 Fusion Encoder
Using the masks directly to train a classification network is not ideal (see Table 4) because the masks lack texture details and are polarized to all-or-nothing errors due to their coarse discretized representation. Hence, we propose a novel Fusion Encoder that uses a one-directional feature fusion scheme (from mask network to image), at the deeper layers (evidence for this design in Section 3.5) to exploit the learned semantics from both networks. The Fusion Encoder consists of an image network, a mask network, fusion blocks, and a classifier.
Image and Mask Network: We use two 5 staged lightweight ShuffleNet [11] backboned encoders as our image and mask network. In the first three stages, both networks learn the features independent from each other. After the third stage, features from the image and mask networks are fused and passed to the fourth stage of the image network. The features from the two networks are fused again before the fifth stage and passed to the image network to obtain the final features from the image network.
Fusion Blocks and Classifier Fusion blocks are placed at the beginning of the $4^{th}$ $5^{th}$ stages of the networks and consist of a pixel-wise summation of the image and mask features, followed by a convolution, batch normalization, and ReLU operations. A pixel-wise summation is performed between the final features obtained at the end of the image and mask networks, followed by average pooling and fed through a randomly initialized fully connected linear classifier that outputs (after softmax & argmax) the morphology class predictions of the sperm head crop.
@@ 2.3 Soft Mixup Regularization
Accurately classifying the sperm head morphology is difficult even for experts, which may lead to disagreement on expert-labeled class IDs of samples (i.e., for SCIAN a sample may have a majority class label with consensus among at least 2/3 expert labels and a minority class label with no consensus).
Soft Mixup Augmentation: We apply an intra-class Mixup augmentation [20] scheme to handle samples with disagreed labels and reduce class conditional input variance. Given a sample image crop $x_{i}$ and its corresponding mask $m_{i}$ with a majority label $y^{a}_{i}$ and a minority label $y^{b}_{i}$, we randomly pick another sample $x_{j}$ (with a mask $m_{j}$, a majority label $y^{a}_{j}$, and a minority label $y^{b}_{j}$) if $y^{a}_{i}=y^{a}_{j}$. Prior to this step, we perform random oversampling of all the under-represented classes to match the number of samples in the most represented class. We then apply separate image augmentation ($\Phi$) to $x_{i}$ and $x_{j}$. The input, $\ddot{x}_{i}$, to the Fusion Encoder is a linear combination of the augmented mixup samples $\Phi_{i}(x_{i})$ and $\Phi_{j}(x_{j})$, and its target label $\ddot{y}_{i}$ is equivalent to the majority label ($y_{i}^{a}$). Formally:
@@
Soft Mixup Loss: We modify the Soft Loss in [21] to adapt intra-class mixup to handle partial agreements among the morphology class labels. Soft Loss is defined as:
@@ 3.1 Datasets
We demonstrate the performance of SHMC-Net on two public sperm morphology datasets. SCIAN-Morpho Sperm GS (SCIAN) [1] contains 1854 gray-scale sperm images of size 35 $\times$ 35 each belonging to one of five morphological classes: Normal (N), Tapered (T), Pyriform (P), Small (S), or Amorphous (A) as labeled by 3 experts. Samples where all 3 experts agree on the same class label form the SCIAN total Agreement (SCIAN-TA) dataset (384 images: 35 N, 69 T, 7 P, 11 S, 262 A), and the samples where at least 2 experts agree on class labels form the SCIAN Partial Agreement (SCIAN-PA) dataset (1132 images: 100 N, 228 T, 76 P, 72 S, 656 A)
@@ 3.2 Implementation Details
ImageNet pre-trained ShuffleNet V2 [11] is used as the backbone for our image and mask networks and trained using the AdamW optimizer with an initial learning rate of 0.00015 and the Cosine Annealing learning rate scheduler. The model is trained on a single NVIDIA TITAN-Xp GPU for 300 epochs (batch size of 128 for SCIAN-PA and 64 for SCIAN-TA and HuSHeM). Dropout with 40% probability and weight decay of 0.01 was applied as additional regularization. The majority class weight $\gamma$ is set to 0.85 and mixup strength, $\lambda$, is sampled randomly from the beta distribution with $\alpha=0.5$. All the image crops are first resized to 64$\times$64 and randomly square-cropped between sizes 45 to 64 and resized again to 64*64 along with 10-degree rotation, vertical flip, and horizontal/vertical shift by up to 10% of their width or height. We evaluate the performance in terms of Overall Accuracy, macro-averaged class-wise Precision and Recall, and the F1 score attained as a harmonic mean of the two. We apply 5-fold cross-validation and report the average of 5 runs.
@@ 3.5 Component Studies
Method Accuracy
Precision
Recall F1
@@
Spencer et al. [18] 98.52
98.53
98.52 98.52
SHMC-Net-E (ours) 99.17 $\pm$ 0.20
99.19 $\pm$ 0.21
99.16 $\pm$ 0.20 99.17 $\pm$ 0.20
@@ 3.6 Ablation Study
Combining mask and image features at the classifier level has improvements (3.4% F1 on SCIAN-PA and 1.8% on HuSHeM) over using image-only. Intermediate stage fusion of mask and image features further strengthens the model’s ability to learn morphology features (further 1.9% on SCIAN-PA and 1.3% on HuSHeM). Finally, handling class imbalance and label discrepancies through Soft Mixup adds 2.2% further F1 improvement on SCIAN-PA, and the regularization through Soft Mixup contributes to an additional 0.4% F1 on the small-size HuSHeM dataset.